### v2

@@ Introduction
The present review summarizes the state of our understanding, in the third decade of the 21st century, of the nuclear reactions and decays taking place in solar and stellar hydrogen burning. It also addresses related issues, including solar neutrino detection, radiative opacities, electron screening of nuclear reactions, and the status of the current facilities for measuring cross sections and opacities.
As was the case for two previous decadal reviews [1998RvMP...70.1265A, 2011RvMP...83..195A], the present review summarizes the progress made over the past decade in advancing the nuclear physics of main-sequence stars and makes recommendations for future work.
@@ I.1 Purpose
Ray Davis’s measurements of the flux of solar neutrinos [1968PhRvL..20.1205D, 2003RvMP...75..985D] showed a significant deficit with respect to the predictions of the standard solar model (SSM) [1966PhRvL..17..398B]. This discrepancy became known as the "solar neutrino problem." An intense debate ensued about possible explanations for this discrepancy, summarized in the entertaining review by 1988Natur.334..487B. Suggestions included plausible flaws in the SSM that might produce a cooler solar core (e.g. 1981A&A....96....1S; 1996PhRvL..77.4286C), new particle physics such as neutrino oscillations (1969PhLB...28..493G; 1978PhRvD..17.2369W), changes in the nuclear physics governing He synthesis, such as an unidentified low-energy resonance in the 3He+3He reaction affecting extrapolations of laboratory cross sections to solar energies [1972Natur.238...24F, 1972PhLB...40..602F], or the presence of dark matter particles modifying energy transport in the solar core [1985ApJ...294..663S]. By the mid-1990s, with new results from Kamioka II/III (1989PhRvL..63...16H; 1996PhRvL..77.1683F) and the SAGE/GALLEX (1991PhRvL..67.3332A; 1992PhLB..285..376A) experiments revealed a pattern of neutrino fluxes inconsistent with the expected scaling of those fluxes with the solar core temperature. Attention increasingly turned to neutrino oscillations and other potential particle physics solutions. Concerns about the nuclear physics of the SSM also evolved, focusing more on the uncertainties that might inhibit extraction of any such “new physics.”
@@
At that time, during a meeting on the Solar Neutrino Problem hosted by the Institute for Nuclear Theory (INT), University of Washington, Seattle, a suggestion was made to convene the community working on solar nuclear reactions, in order to reach consensus on the best current values of cross sections and their uncertainties. The INT hosted the proposed workshop in February, 1997, drawing representatives from almost every experimental group active in this area, as well as many of the theorists who were engaged in solar neutrino physics. The working groups that formed during this meeting worked over the following year to evaluate past work, determining the needed cross sections and their uncertainties. The results of this evaluation, which came to be known as Solar Fusion I or 1998RvMP...70.1265A [1998RvMP...70.1265A], became the standard for use in solar modeling over the following decade.
@@
An update of 1998RvMP...70.1265A was launched with the January, 2009, workshop “Solar Fusion Cross Sections II”, hosted again by the INT in Seattle. Initial results from SuperKamiokande, the Sudbury Neutrino Observatory, and Borexino were then in hand, and the conversion of approximately two thirds of solar electron neutrinos into other flavors had been firmly established [2002PhRvL..89a1301A]. Consequently, the motivation for the study was to ensure that solar model predictions would be based on the most current nuclear physics, so that meaningful uncertainties could be placed on neutrino parameters derived from solar neutrino measurements, such as the mixing angle . An additional source of uncertainty had arisen at the time, with the advent of three-dimensional (3D) radiative-hydrodynamic models of the Sun’s atmosphere for determining element abundances from photo-absorption lines [2006NuPhA.777....1A, 2009ARA&A..47..481A]. This improved analysis, though, lowered the inferred metallicities, leading to tension with the Sun’s interior sound speed profile determined from helioseismology [2002RvMP...74.1073C]. The discrepancy between higher metallicity models that accurately reproduce the Sun’s interior sound speed and lower metallicity models based on the most current treatment of the photosphere was termed the “solar composition problem”. While the neutrino fluxes were not known to the precision needed to distinguish between the competing models, it had been shown that the solar core metallicity could be extracted directly from future CN solar neutrino measurements, if the precision of associated nuclear cross sections were improved [2008ApJ...687..678H]. Reflecting the impact of new experimental work and improved theory, the 2011RvMP...83..195A study of 2011RvMP...83..195A significantly revised the 1998RvMP...70.1265A cross section for the driving reaction of the CNO-I cycle, and updated key pp-chain cross sections.
@@
The 47 workshop participants represented most of the leading experimental and theoretical research groups that are active in the field. Workshop participants were organized into nine working groups, who were charged with reviewing and evaluating work completed since 2011RvMP...83..195A, to produce updated recommendations. The structure of this review largely follows that of the working groups. The timing of SF III was driven in part by planned large-scale asteroseismic surveys of hydrogen-burning stars such as ESA’s PLATO333https://platomission.com/ mission [2014ExA....38..249R], and the expectation that associated data analysis will require large libraries of stellar models in which masses, ages, metallicities, and other parameters are varied. SF III will provide the best current nuclear astrophysics input for such modeling. This review summarizes that input, based on the recommendations of the nine SF III working groups.
@@ I.2 Terminology used
The terminology used here follows recent practice, as summarized for example in the NACRE compilation [1999NuPhA.656....3A]. It is briefly described to aid the reader.
@@
with and the charge numbers of the target and projectile, the fine structure constant, the relative velocity of the interacting particles in units of , the reduced mass, and the center-of-mass energy. (In this review we use energy units of MeV unless otherwise stated.) This removes the sharp energy dependence from the cross section associated with s-wave Coulomb scattering off a point nucleus, thereby accounting for the leading penetration effects associated with the Coulomb barrier. The quantity is known as the Sommerfeld parameter. Consequently, unless there are resonances, varies gently, and at low energies can generally be expressed as a low-order polynomial in .
The S-factor in Eq. (2) is defined in terms of the nuclear (or “bare") cross section, while laboratory experiments are performed with targets consisting of neutral atoms. As electrons shield the nucleus, lowering the Coulomb barrier, their presence enhances the laboratory cross section. The S-factors derived from laboratory data must be corrected for this effect. As the energy is lowered, the screening correction becomes increasingly important. Thus particular care must be taken when extrapolating laboratory cross sections to lower energies, as Fig. 4 of 2011RvMP...83..195A illustrates. Unless otherwise noted, the S-factors discussed here are those for bare nuclei.

### v3

@@ Abstract
- I Introduction
- II Solar neutrino observations
- III The 1HH reaction ()
- IV The reaction ()
- V The reaction )
- VI The reaction ()
- VII The hep reaction ()
- VIII Electron capture by p+p and
- IX The reaction ()
- X The reaction (S1 14)
- XI Other CNO, Ne-Na reactions
- XII Electron screening of nuclear reactions
- XIII Radiative opacities
- XIV Experimental facilities for solar fusion studies
- XV Closing remarks
- A Gallium Neutrino Source Cross Sections
- B Bayesian Methods
@@ Introduction
The present review summarizes the state of our understanding, in the third decade of the 21st century, of the nuclear reactions and decays taking place in solar and stellar hydrogen burning. It also addresses related issues, including solar neutrino detection, radiative opacities, electron screening of nuclear reactions, and the status of the current facilities for measuring cross sections and opacities. As was the case for two previous decadal reviews [11, 12], the present review summarizes the progress made over the past decade in advancing the nuclear physics of main-sequence stars and makes recommendations for future work.
@@ I.1 Purpose
Ray Davis’s measurements of the flux of solar neutrinos [225, 224] showed a significant deficit with respect to the predictions of the standard solar model (SSM) [67]. This discrepancy became known as the "solar neutrino problem." An intense debate ensued about possible explanations for this discrepancy, summarized in the entertaining review by Bahcall et al. [75]. Suggestions included plausible flaws in the SSM that might produce a cooler solar core (e.g. Schatzman et al. 573; Cumming and Haxton 207), new particle physics such as neutrino oscillations (Gribov and Pontecorvo 313; Wolfenstein 662), changes in the nuclear physics governing He synthesis, such as an unidentified low-energy resonance in the 3He+3He reaction affecting extrapolations of laboratory cross sections to solar energies [275, 268], or the presence of dark matter particles modifying energy transport in the solar core [602]. By the mid-1990s, with new results from Kamioka II and III (Hirata et al. 353; Fukuda et al. 286) and the SAGE/GALLEX (Abazov et al. 1; Anselmann et al. 47) experiments revealed a pattern of neutrino fluxes inconsistent with the expected scaling of those fluxes with the solar core temperature. Attention increasingly turned to neutrino oscillations and other potential particle physics solutions. Concerns about the nuclear physics of the SSM also evolved, focusing more on the uncertainties that might inhibit extraction of any such “new physics.”
@@
At that time, during a meeting on the Solar Neutrino Problem hosted by the Institute for Nuclear Theory (INT), University of Washington, Seattle, a suggestion was made to convene the community working on solar nuclear reactions, in order to reach consensus on the best current values of cross sections and their uncertainties. The INT hosted the proposed workshop in February, 1997, drawing representatives from almost every experimental group active in this area, as well as many of the theorists who were engaged in solar neutrino physics. The working groups that formed during this meeting worked over the following year to evaluate past work, determining the needed cross sections and their uncertainties. The results of this evaluation, which came to be known as Solar Fusion I or SF I [11], became the standard for use in solar modeling over the following decade.
@@
An update of SF I was launched with the January, 2009, workshop “Solar Fusion Cross Sections II”, hosted again by the INT in Seattle. Initial results from SuperKamiokande, the Sudbury Neutrino Observatory, and Borexino were then in hand, and the conversion of approximately two thirds of solar electron neutrinos into other flavors had been firmly established [23]. Consequently, the motivation for the study was to ensure that solar model predictions would be based on the most current nuclear physics, so that meaningful uncertainties could be placed on neutrino parameters derived from solar neutrino measurements, such as the mixing angle . An additional source of uncertainty had arisen at the time, with the advent of three-dimensional (3D) radiative-hydrodynamic models of the Sun’s atmosphere for determining element abundances from photo-absorption lines [57, 59]. This improved analysis, though, lowered the inferred metallicities, leading to tension with the Sun’s interior sound speed profile determined from helioseismology [193]. The discrepancy between higher metallicity models that accurately reproduce the Sun’s interior sound speed and lower metallicity models based on the most current treatment of the photosphere was termed the “solar composition problem”. While the neutrino fluxes were not known to the precision needed to distinguish between the competing models, it had been shown that the solar core metallicity could be extracted directly from future carbon-nitrogen (CN) solar neutrino measurements, if the precision of associated nuclear cross sections were improved [340]. Reflecting the impact of new experimental work and improved theory, the SF II study of Adelberger et al. [12] significantly revised the SF I cross section for the driving reaction of the CNO-I cycle, and updated key pp-chain cross sections.
@@
The 47 workshop participants represented most of the leading experimental and theoretical research groups that are active in the field. Workshop participants were organized into nine working groups, who were charged with reviewing and evaluating work completed since SF II, to produce updated recommendations. The structure of this review largely follows that of the working groups. The timing of SF III was driven in part by planned large-scale asteroseismic surveys of hydrogen-burning stars such as ESA’s PLATO333https://platomission.com/ mission [543], and the expectation that associated data analysis will require large libraries of stellar models in which masses, ages, metallicities, and other parameters are varied. SF III will provide the best current nuclear astrophysics input for such modeling. This review summarizes that input, based on the recommendations of the nine SF III working groups.
@@ I.2 Terminology used
The terminology used here follows recent practice, as summarized for example in the Nuclear Astrophysics Compilation of Reaction Rates (NACRE) compilation [44]. It is briefly described to aid the reader.
@@
with and the charge numbers of the target and projectile, the fine structure constant, the relative velocity of the interacting particles in units of , the reduced mass, and the center-of-mass energy. (In this review we use energy units of MeV unless otherwise stated.) This removes the sharp energy dependence from the cross section associated with s-wave Coulomb scattering off a point nucleus, thereby accounting for the leading penetration effects associated with the Coulomb barrier. The quantity is known as the Sommerfeld parameter. Consequently, unless there are resonances, should be a smooth function that at low energies can generally be expressed as a low-order polynomial in .
The S-factor in Eq. (2) is defined in terms of the nuclear (or “bare") cross section, while laboratory experiments are performed with targets consisting of neutral atoms. As electrons shield the nucleus, lowering the Coulomb barrier, their presence enhances the laboratory cross section. The S-factors derived from laboratory data must be corrected for this effect. As the energy is lowered, the screening correction becomes increasingly important. Thus particular care must be taken when extrapolating laboratory cross sections to lower energies, as Fig. 4 of SF II illustrates. Unless otherwise noted, the S-factors discussed here are those for bare nuclei.